The table describes hypothetical employment statistics in the United States in a given year. The labor force is the sum of those that are employed plus the unemployed. Labor Full-time workers and UnderemployedDiscouraged force voluntary part-workers workersworkers million5 million 155 140 million million Based on the information, what is the unemployment rate? Specify your answer to one decimal place unemployment rate
Answer #1

Labor force = Employed + unemployed

Employed = (140+7)= 147 million

Labor force = 155 million

Unemployed = (155-147) million = 8 million

Unemployment rate = (8/155) 100 =5.16% = 5.2%